CS621 – Syllabus Algorithm Design and Analysis

Instructor: B. David Saunders

Office: 456 Smith Hall  Hours: 2:30AM-4:00PM TuTh and by arrangement
Phone: 831-6238  Email: saunders@cis.udel.edu

Teaching Assistant: Rangarajan Vijayaraghavan

Office: 115 Pearson Hall  Hours: 2:00-3:00 Tuesday and 11:00-12:00 Thursday
Phone: 831-1949  Email: rangaraj@cis.udel.edu

Textbook

Cormen, Leiserson, and Rivest Introduction to Algorithms, McGraw-Hill & MIT Press, 1990.

Other references. See CLR for title, publisher, date. Aho, Hopcroft, and Ullman: Classic algorithms text – crystalized the topic. Garey and Johnson: famous summary of NP Completeness, enumeration of many NP complete problems. Sedgewick: Well done text, interesting algorith choice, well explained, weak on analysis. Purdom and Brown: Unusual treatment, good source on mathematical methods.

1 Coverage and pace

We will cover most of the algorithms discussed in CLR, roughly in the order given. There will be some skipping and there may be some topics not from the text. At each Lecture I will indicate the reading appropriate for the topic of the next lecture. After each lecture I will update an online record of lecture topics and the related readings. Any sections being skipped and not relevant to any of the homework will be indicated also.

Lecture 1: The first day we used as an example the Karatsuba integer multiplication algorithm used in the multiprecision arithmetic package from the gnu library. Not in the text. First reading: Scan chapters 1-6 for future reference, pay closer attention to 2 and 4 which are immediately applicable. Next up: Topics on sorting Chapters 7, 8, and 9.1

2 Grading

80% Homeworks (5 or 6 homework problem sets comprising 25-30 problems in total) 20% Final Exam in exam week (factual questions about the algorithms) fudge factor: classroom participation and other evidence of performance.

Policy on plagarism: All homework in this course is designed to be done individually. The giving or taking of another person’s work (with or without modification of detail) is plagarism and will be handled in accordance with University procedures.